Grade 11Mechanics

THE RATE OF DOING WORK BY FORCE ACTING ON ABODY MOVINFG ALONG X-AXIS DEPENDS ON POSITION X OF BODY AND IS EQUAL TO 2X . THE VELOCITY OF PARTICLE IS GIVEN BY EXPRESSION

JODU SHANKAR SINGH
8 Years agoGrade 11

1 Answer

Hari Sundar Pyda
8 Years ago
Position(or displacement)= x
Rate of doing work= power= 2x
Let the time interval be t. 
Power= work/time
2x= force*displacement/t
2x*t= F*x
2t= F
2t=ma
2t=mv/t
v=2t2/m= velocity.
Hope it helps!!